# Neural Radiance Fields in Medical Imaging: Challenges and Next Steps

### 26 Feb 2024 v1

@@ 3 Methods
Fig. 3: Over view of Uncertainty-aware MedNeRF [6]. Top: Network structures and training phrase [5]. Mid: Fine-tuning the generator for balancing the Blurrinenss and Accuracy. Bottom: Rendering the whole 3D volume (each view taken at a 5-degree interval) using a single view slide.
@@ 3.1 NeRFs for Different Organs
Fig. 5: Overview of Masked NeRF [13] Left: Skull projections from NeRF. Right. Skull projections from Masked NeRF scene representations.
@@ 4.1 Overview
Training Medical Image NeRFs needs datasets comprising paired 2D (e.g. X-ray) and 3D (e.g. CT, MRI) scans. However, assembling a real dataset of such pairs is virtually impractical due to the rarity of obtaining X-ray pairs and CT (or MRI) scans captured simultaneously in clinical practice [22]. Thus, the Reconstructed Radiographs (DRR) technique is employed for simulating and generating X-ray images. This method entails the creation of virtual X-ray projections using three-dimensional medical image data, such as CT or MRI. By generating DRRs, we can avoid the need of paired X-rays and corresponding CT reconstructions, which would otherwise subject patients to increased radiation exposure. The overview relation between Digitally Reconstructed Radiograph (DRR) and medical imaging NeRFs is shown in Figure 8.
@@ 4.2 Digitally Reconstructed Radiographs (DRR)
Fig. 8: The relation between Digitally Reconstructed Radiograph (DRR) and medical imaging NeRFs [23]. Left: leverages original CT scan slices, Mid: perceived as a three-dimensional voxel array, Right: to generate a DRR. This direct volume rendering scheme efficiently transforms the volumetric data from a CT scan into a visually comprehensive DRR, offering a synthesized view similar to traditional X-ray imagery.
@@ 4.3 Public Dataset
PerX2CT [22] (Github [30]). To reconstruct CT images from X-rays, a dataset comprising paired X-ray and CT scans is essential. However, assembling a real dataset of such pairs is virtually impractical due to the rarity of obtaining X-ray pairs and CT scans captured simultaneously in clinical practice. Consequently, existing research has turned to digitally reconstructed radiograph (DRR) technology as a solution, enabling the generation of synthetic X-rays from actual CT volumes to simulate the required paired data for studies, the dataset is constructed following [31].
@@ 4.4 Evaluation Metrics
2D Rendering. Another evaluation of the model focuses on the task of 2D rendering, utilizing the Frechet Inception Distance (FID) and Kernel Inception Distance (KID) metrics for a comprehensive comparison of image quality, where lower scores signify superior quality [5].
When benchmarking against prevalent 2D rendering methods such as pixelNeRF [32] and GIRAFFE [33, 34] on the datasets, these methods fall short, failing to reach the performance levels of the GRAF baseline or the model. Conversely, the models [5, 6] demonstrate superior performance, outperforming baseline models in terms of both FID and KID metrics across all tested datasets.

### 02 Mar 2024 v2

@@ 3 Methods
Fig. 3: Over view of Uncertainty-aware MedNeRF [6]. Top: Network structures and training phrase [5]. Mid: Fine-tuning the generator for balancing the blurriness and accuracy with uncertainty-aware multi-task loss. Bottom: Rendering the whole 3D volume (each view taken at a 5-degree interval) using a single view slide.
@@ 3.1 NeRFs for Different Organs
Fig. 5: Overview of Masked NeRF [13]. Left: Skull projections from NeRF. Right. Skull projections from Masked NeRF scene representations.
@@ 4.1 Overview
Training Medical Image NeRFs needs datasets comprising paired 2D (e.g. X-ray) and 3D (e.g. CT, MRI) scans. However, assembling a real dataset of such pairs is virtually impractical due to the rarity of obtaining X-ray pairs and CT (or MRI) scans captured simultaneously in clinical practice [22]. For example, gathering paired X-ray and CT data can introduce inaccuracies due to patient movement and variations in equipment. Additionally, this process may expose patients to increased radiation levels [12].
Thus, the Reconstructed Radiographs (DRR) technique is employed for simulating and generating X-ray images [5]. This method entails the creation of virtual X-ray projections using three-dimensional medical image data, such as CT or MRI. By generating DRRs, we can avoid the need for paired X-rays and corresponding CT reconstructions, which would otherwise subject patients to increased radiation exposure. The overview relation between Digitally Reconstructed Radiographs (DRR) and medical imaging NeRFs is shown in Figure 8.
@@ 4.2 Digitally Reconstructed Radiographs (DRR)
Fig. 8: The relation between Digitally Reconstructed Radiograph (DRR) and medical imaging NeRFs, the DRR example is from [23]. Left: leverages original CT scan slices, Mid: perceived as a three-dimensional voxel array, Right: to generate a DRR. This direct volume rendering scheme efficiently transforms the volumetric data from a CT scan into a visually comprehensive DRR, offering a synthesized view similar to traditional X-ray imagery.
@@ 4.3 Public Dataset
PerX2CT Dataset [22] (Github [30]). To reconstruct CT images from X-rays, a dataset comprising paired X-ray and CT scans is essential. However, assembling a real dataset of such pairs is virtually impractical due to the rarity of obtaining X-ray pairs and CT scans captured simultaneously in clinical practice. Consequently, existing research has turned to DRR technology as a solution, enabling the generation of synthetic X-rays from actual CT volumes to simulate the required paired data for studies, the dataset is constructed following [31].
DIF-Net Dataset [32] (Github [33]). The dataset comprises a knee CBCT collection with 614 CT scans, distributed as 464 for training, 50 for validation, and 100 for testing purposes. It also includes 2D projections created using DRR technology, each with a resolution of 256 $\times$× 256.
@@ 4.4 Evaluation Metrics
2D Rendering. Another evaluation of the model focuses on the task of 2D rendering, utilizing the Frechet Inception Distance (FID) and Kernel Inception Distance (KID) metrics for a comparison of image quality, where lower scores signify superior quality [5]. For example, the models [5, 6] demonstrate superior performance, outperforming baseline models, such as pixelNeRF [34] and GIRAFFE [35, 36], in terms of both FID and KID metrics across all tested datasets.